How to translate into equations to be solved? Thanks!
In a triangle, the sum of the interior angles is 180 degrees. If angle A is three times as large as angle B, and angles B and C sum to 90 degrees, what is the measure of each angle?

Answer by snovember85(4)

Since the sum of the interior angles is 180, then A+B+C=180. A is 3 times larger than B, so A = 3B. Notice that B+C=90, that means A = 90, because the sum of three angles is 180. So (A =) 3B = 90, then B = 90/3 = 30. Knowing B is 30, you can figure out C = 90 - 30 = 60.